Soil composition of a loam that mixed with water in a jar and left to stand for 3 day..1.1.2 measure the thickness of the sand ;clay and slit layers in your jar. a)sand layer thichness?

b)clay layer thickness?

C) slit layer thickness?



1
Expert's answer
2021-09-14T01:53:58-0400

The question states;

1:1:2 for sand, clay and silt respectively


"\%" sand = 1/4 × 100% = 25%

"\%" clay = 2/4 × 100% = 50%

"\%" silt = 1/4 × 100% = 25%